# Knowledge Editing for Large Language Models: A Survey

### v3

@@ 7.3. Natural Language Generation
7.3.1. Background.
KME techniques are also promising to ensure the relevancy of the Natural Language Generation (NLG) task, which aims to generate coherent and contextually relevant content based on provided instructions (REITER and DALE, 1997). Considering the rapid evolution of the global information landscape, it is essential for NLG models to remain up-to-date and ensure the accuracy of generated text while avoiding potentially false statements that may mislead the users.
7.3.2. Existing Works.
In practice, several works have been proposed to apply KME methods to promote model performance in natural language generation tasks. For instance, FRUIT (au2 et al., 2022) proposes to update outdated Wikipedia articles according to the collection of new information about the article’s subject. Based on the T5 model (Raffel et al., 2020), the authors utilize a compressed output format to eliminate the necessity of generating the entire update from scratch and promote thoughtful content structuring, which effectively handles the challenge of incoherence. In MEND (Mitchell et al., 2022a), the authors apply their proposed method in the Wikitext generation task, where the edited model is required to produce credible 10-token extensions based on a provided Wikitext prefix (Ma, 2021). With modification on multi-layer token-wise activations and gradients, the edited model presents higher coherence on the NLG task, which demonstrates the effectiveness of KME in generating target texts with richer information than QA or FC.
@@ 8. Discussion
Despite the continual progress of works on KME, several critical aspects have been inadequately addressed by existing studies. Delving deeper into these challenges could offer researchers fresh insights and pave the way for the further advancement of the field. Consequently, in this section, we outline the pressing challenges that await solutions in KME.
@@ 8.1. Challenges
8.1.1. Trade-off between Locality and Generality.
@@
In KME, it is crucial to balance two objectives, locality and generality (as defined in Sec. 4), such that a higher edit success rate can be achieved with minimal negative influence on other knowledge irrelevant to the edits. When editing a language model, a potential trade-off might emerge between these two desirable properties. As demonstrated in (Yao et al., 2023), local modification methods, such as MEMIT (Meng et al., 2023) and ROME (Meng et al., 2022) generally preserve a higher level of locality, as they locate precise locations of target knowledge to conduct the edition, which does not largely affect the unrelated weights. In addition, T-Patcher (Huang et al., 2023) points out that increasing the size of memory increases locality while decreasing the generality. These observations underscore the intricate balance between locality and generality. However, it remains challenging to tackle the trade-off problem and achieve a balance between these two desirable properties of KME methods.
8.1.2. Theoretical Analysis.
@@
While many current KME studies focus on developing effective methods to enhance the editing performance regarding various desirable properties, there exists a notable gap between the practical application and the comparatively less discovered theoretical analysis. Recently, in (Tanno et al., 2022), the authors provide theoretical support for the justification of identifying harmful training examples and editing the model by erasing the information from a Bayesian view. LEACE (Belrose et al., 2023) introduces an analytical framework that offers a theoretical perspective for the task of erasing target concept information from every layer in language models. In general, the benefits of incorporating theoretical analysis are multi-faceted. First, theoretical analysis provides a deeper understanding of the mechanics underlying KME, allowing for more principled approaches to editing. Second, a strong theoretical basis sets a solid foundation for future research, encouraging more rigorous and systematic exploration in the field of KME. However, to the best of our knowledge, there still does not exist any comprehensive theoretical analysis regarding the KME problem that involves novel knowledge. We hope that future research will enrich the theoretical discourse that can deliver profound insights into the substantial foundations of KME methods.
8.1.3. Editing at Scale.
@@
Another crucial property that hinders the practical application of KME is scalability — the ability of editing strategy to effectively perform a large number of edits simultaneously (Mitchell et al., 2022a). For example, conversational systems (Zheng et al., 2023) are expected to be constantly updated to incorporate an enormous number of global events and the information originating from them. However, as the number of applied edits increases, the coherence of language models is severely jeopardized, as multiple edits might contradict a broader spectrum of pre-existing knowledge in the models (Wang et al., 2023b). This can lead to decreased editing performance in both locality and generality metrics (Mitchell et al., 2022b). Although external memorization methods can alleviate such problems with a larger size of memories of additional parameters, they are still vulnerable if thousands of edits are required (Meng et al., 2022). Moreover, simply adapting single-edit techniques for a multi-edit environment by merely applying them sequentially has been demonstrated to be proven suboptimal (Meng et al., 2023). Therefore, the unique and intricate challenge of coherence renders editing at scale a formidable task.
@@ 8.2. Future Directions
Auto-Discovery of Editing Targets. Current KME methods mainly rely on human expertise to identify and incorporate desirable knowledge into pre-trained LLMs. This approach is inherently labor-intensive and can incur significant costs, especially considering the vast and rapidly expanding new information needed to be integrated into language models. A promising future direction lies in the automation of the edits, which aims to identify, evaluate, and prioritize new knowledge that needs to be integrated from raw resources such as websites and social media. Through this strategy, the application of KME can be streamlined, rendering it more practical and adaptable in real-world scenarios. A straightforward solution would be crawling new knowledge and transforming it into a knowledge base, querying LLMs for each knowledge triple, and editing the wrong answer. However, such a strategy still lacks efficiency. Therefore, it remains a crucial task to discover editing knowledge from various resources without human effort.
@@
Continual Editing. Current methodologies on LLM editing primarily consider one-step offline editing (De Cao et al., 2021; au2 et al., 2022); however, such an approach is not aligned with real-world applications where models might continually encounter novel knowledge to be injected. For example, an online question-answering (QA) model may continually encounter reports of incorrect answers from end users, where the editing needs to be conducted on the run (Huang et al., 2023). Therefore, an optimal KME technique should be capable of instantaneously and continuously rectifying emergent issues. We note that continual editing of pre-trained LLMs presents a unique challenge: preventing the edited models from forgetting or contradicting previous edits. Despite the inherent complexities, the persistent demand for continual editing in practice underscores the importance of solving this challenge.
@@
Robust Editing. An important direction for the advancement of KME lies in enhancing its robustness. In an era where misinformation spreads rapidly, it is urgent that edited models not only retain their accuracy but also resist adversarial attacks and misinformation (Ganguli et al., 2022). Here, we should note that the concept of robustness extends beyond just maintaining factual accuracy; it involves fortifying the model against potentially adversarial external perturbations (Perez et al., 2022). For example, if KME is maliciously applied to inject harmful knowledge into language models, the edited models can be easily transformed into tools for misinformation (Taori et al., 2023). Therefore, to prevent such cases, it is crucial for KME techniques to develop capabilities that can identify and counteract such unwanted inputs, thereby enhancing their resilience against adversarial actions. In practice, as the trend leans towards open-sourcing large language models, it becomes ever more crucial to safeguard against potential manipulations that can turn these models harmful.
@@
Editable Fairness. With the wide application of large language models (LLMs) to support decisions, the emphasis on fairness has grown significantly (Wang et al., 2023a), which requires LLMs to fairly treat people with diverse background (Abid et al., 2021). However, LLMs trained on large datasets inevitably incorporate certain biases during this pre-training phase  (Dong et al., 2019). Fortunately, the precision and efficiency of KME techniques offer a promising solution to mitigate such biases and promote fairness in pre-trained LLMs. For instance, in a model designed to classify biographical sentences with occupation (De-Arteaga et al., 2019), KME can be used to inject nuanced knowledge about a particular profession, guiding the model towards a more equitable understanding of individuals associated with that profession (Hernandez et al., 2023). However, this remains a complex challenge, as fairness often entails considering disparate groups of individuals rather than specific people. This broader focus makes knowledge injection via KME a non-trivial task. Despite these difficulties, the enhancement of fairness in language models is paramount, and KME techniques present a promising avenue to achieve this goal.

### v4

@@ 7.3. Natural Language Generation
Background. KME techniques are also promising to ensure the relevancy of the Natural Language Generation (NLG) task, which aims to generate coherent and contextually relevant content based on provided instructions (REITER and DALE, 1997). Considering the rapid evolution of the global information landscape, it is essential for NLG models to remain up-to-date and ensure the accuracy of generated text while avoiding potentially false statements that may mislead the users.
Existing Works. In practice, several works have been proposed to apply KME methods to promote model performance in natural language generation tasks. For instance, FRUIT (au2 et al., 2022) proposes to update outdated Wikipedia articles according to the collection of new information about the article’s subject. Based on the T5 model (Raffel et al., 2020), the authors utilize a compressed output format to eliminate the necessity of generating the entire update from scratch and promote thoughtful content structuring, which effectively handles the challenge of incoherence. In MEND (Mitchell et al., 2022a), the authors apply their proposed method in the Wikitext generation task, where the edited model is required to produce credible 10-token extensions based on a provided Wikitext prefix (Ma, 2021). With modification on multi-layer token-wise activations and gradients, the edited model presents higher coherence on the NLG task, which demonstrates the effectiveness of KME in generating target texts with richer information than QA or FC.
@@ 8.1. Challenges
Despite the continual progress of works on KME, several critical aspects have been inadequately addressed by existing studies. Delving deeper into these challenges could offer researchers fresh insights and pave the way for the further advancement of the field. Consequently, we hereby outline the pressing challenges that await solutions in KME.
@@
Trade-off between Locality and Generality. In KME, it is crucial to balance two objectives, locality and generality (as defined in Sec. 4), such that a higher edit success rate can be achieved with minimal negative influence on knowledge irrelevant to the edits. When editing a language model, a potential trade-off might emerge between these two desirable properties. As demonstrated in (Yao et al., 2023), local modification methods, such as MEMIT (Meng et al., 2023) and ROME (Meng et al., 2022) generally preserve a higher level of locality, as they locate precise locations of target knowledge to conduct the edition, which does not largely affect the unrelated weights. In addition, T-Patcher (Huang et al., 2023) points out that increasing the size of memory increases locality while decreasing the generality. These observations underscore the intricate balance between locality and generality. However, it remains challenging to tackle the trade-off problem and achieve a balance between these two desirable properties of KME methods.
@@
Theoretical Analysis. While many current KME studies focus on developing effective methods to enhance the editing performance regarding various desirable properties, there exists a notable gap between the practical application and the comparatively less discovered theoretical analysis. Recently, in (Tanno et al., 2022), the authors provide theoretical support for the justification of identifying harmful training examples and editing the model by erasing the information from a Bayesian view. LEACE (Belrose et al., 2023) introduces an analytical framework that offers a theoretical perspective for the task of erasing target concept information from every layer in language models. In general, the benefits of incorporating theoretical analysis are multi-faceted. First, theoretical analysis provides a deeper understanding of the mechanics underlying KME, allowing for more principled approaches to editing. Second, a strong theoretical basis sets a solid foundation for future research, encouraging more rigorous and systematic exploration in the field of KME. However, to the best of our knowledge, there still does not exist any comprehensive theoretical analysis regarding the KME problem that involves novel knowledge. We hope that future research will enrich the theoretical discourse that can deliver profound insights into the substantial foundations of KME methods.
@@
Editing at Scale. Another crucial property that hinders the practical application of KME is scalability — the ability of editing strategy to effectively perform a large number of edits simultaneously (Mitchell et al., 2022a). For example, conversational systems (Zheng et al., 2023) are expected to be constantly updated to incorporate an enormous number of global events and the information originating from them. However, as the number of applied edits increases, the coherence of language models is severely jeopardized, as multiple edits might contradict a broader spectrum of pre-existing knowledge in the models (Wang et al., 2023c). This can lead to decreased editing performance in both locality and generality metrics (Mitchell et al., 2022b). Although external memorization methods can alleviate such problems with a larger size of memories of additional parameters, they are still vulnerable if thousands of edits are required (Meng et al., 2022). Moreover, simply adapting single-edit techniques for a multi-edit environment by merely applying them sequentially has been demonstrated to be proven suboptimal (Meng et al., 2023). Therefore, the unique and intricate challenge of coherence renders editing at scale a formidable task.
Unstructured Editing. KME faces significant challenges due to its evaluation strategies that focus on knowledge triples, e.g., $t=(s,r,o)$, which are not reflective of how real-world knowledge updates occur (Zhang et al., 2024; Huang et al., 2024). In reality, updates are often found in unstructured texts such as news articles and scientific papers. To address this gap, a recent benchmark (Wu et al., 2024), namely UKE (Unstructured Knowledge Editing), is proposed to evaluate editing performance using unstructured texts as knowledge updates. The experimental results demonstrate significant performance declines of state-of-the-art KME methods. Notably, such a decline persists even with knowledge triplets extracted from unstructured texts. As such, it is imperative to develop more robust and adaptable methods that use unstructured texts for editing.
@@ 8.2. Future Directions
Auto-Discovery of Editing Targets. Current KME methods mainly rely on human expertise to identify and incorporate desirable knowledge into pre-trained LLMs (Yao et al., 2023; Wu et al., 2024; Zhang et al., 2024). This approach is inherently labor-intensive and can incur significant costs, especially considering the vast and rapidly expanding new information needed to be integrated into language models. A promising future direction lies in the automation of the edits, which aims to identify, evaluate, and prioritize new knowledge that needs to be integrated from raw resources such as websites and social media. Through this strategy, the application of KME can be streamlined, rendering it more practical and adaptable in real-world scenarios. A straightforward solution would be crawling new knowledge and transforming it into a knowledge base, querying LLMs for each knowledge triple, and editing the wrong answer. However, such a strategy still lacks efficiency. Therefore, it remains a crucial task to discover editing knowledge from various resources without human effort.
@@
Continual Editing. Current KME methods primarily consider one-step offline editing (De Cao et al., 2021; au2 et al., 2022); however, such an approach is not aligned with real-world applications where models might continually encounter novel knowledge to be injected. For example, an online question-answering (QA) model may continually encounter reports of incorrect answers from end users, where the editing needs to be conducted on the run (Huang et al., 2023). Therefore, an optimal KME technique should be capable of instantaneously and continuously rectifying emergent issues. We note that continual editing of pre-trained LLMs presents a unique challenge: preventing the edited models from forgetting or contradicting previous edits. Despite the inherent complexities, the persistent demand for continual editing in practice underscores the importance of solving this challenge.
@@
Robust Editing. An important direction for the advancement of KME lies in enhancing its robustness. In an era where misinformation spreads rapidly, it is urgent that edited models not only retain their accuracy but also resist adversarial attacks and misinformation (Ganguli et al., 2022). Here, we should note that the concept of robustness extends beyond just maintaining factual accuracy; it involves fortifying the model against potentially adversarial external perturbations (Perez et al., 2022). For example, if KME is maliciously applied to inject harmful knowledge into language models, the edited models can be easily transformed into tools for misinformation (Taori et al., 2023). Therefore, to prevent such cases, it is crucial for KME techniques to develop capabilities that can identify and counteract such unwanted inputs, thereby enhancing their resilience against adversarial actions. In practice, as the trend leans towards open-sourcing LLMs, it becomes ever more crucial to safeguard against potential manipulations that can turn these models harmful.
@@
Editable Fairness. With the wide application of large language models (LLMs) to support decisions, the emphasis on fairness has grown significantly (Wang et al., 2023b), which requires LLMs to fairly treat people with diverse background (Abid et al., 2021). However, LLMs trained on large datasets inevitably incorporate certain biases during this pre-training phase  (Dong et al., 2019). Fortunately, the precision and efficiency of KME techniques offer a promising solution to mitigate such biases and promote fairness in pre-trained LLMs. For instance, in a model designed to classify biographical sentences with occupation (De-Arteaga et al., 2019), KME can be used to inject nuanced knowledge about a particular profession, guiding the model towards a more equitable understanding of individuals associated with that profession (Hernandez et al., 2023). However, this remains a complex challenge, as fairness often entails considering disparate groups of individuals rather than specific people. This broader focus makes knowledge injection via KME a non-trivial task. Despite these difficulties, the enhancement of fairness in language models is paramount, and KME techniques present a promising avenue to achieve this goal.